OpenStudy (littlebird):

Find the derivative of y=cosx/(1+sinx)

OpenStudy (anonymous):

Quotient rule would work. What's stopping you?

OpenStudy (littlebird):

I'm making a mistake somewhere and I can't tell what I'm doing wrong.

OpenStudy (anonymous):

\[ \left(\frac fg \right)'=\frac{f'g-fg'}{g^2} \]

OpenStudy (anonymous):

I need to see work to find a mistake.

OpenStudy (littlebird):

In the book the answer is 1/(1+sinx)

OpenStudy (anonymous):

\[ \sin^2(x)+\cos^2(x)=1 \]

OpenStudy (littlebird):

That's all I needed! Thanks!